OpenStudy (anonymous):

When the quadratic equation y = (x − 4)(2x + 5) is written in standard form, what is the value of the coefficient "b"?

hartnn (hartnn):

do u know what the standard form is?

OpenStudy (anonymous):

Not really ..

hartnn (hartnn):

ok,here it is: \[ax ^{2}+bx+c=0\] can u try to put your given expression into this form??

hartnn (hartnn):

its should be =y ...not =0

OpenStudy (anonymous):

Let me try

OpenStudy (anonymous):

I still haven't figured it out. Can you help me step by step ?

OpenStudy (anonymous):

a. 2 b. 3 c. −3 d. −20

hartnn (hartnn):

sure... y=(x-4)(2x+5) y=x*2x+5x-4*2x-5*4 y=2x^2+5x-8x-20 y=2x^2-3x-20 now compare this with y=ax^2+bx+c what u get as b??

OpenStudy (anonymous):

-3 right ?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

absolutely :)

OpenStudy (anonymous):

Yay, thanks !!